OpenStudy (anonymous):

Calculate the surface area and volume of a spherical ball that has a 2 inch diameter

OpenStudy (anonymous):

SA= 4 pi r^2 V= 4/3 pi r^3

OpenStudy (anonymous):

what s the radius though

OpenStudy (anonymous):

radius is half of you diameter

OpenStudy (anonymous):

1

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so is the answer sa=157.75

OpenStudy (anonymous):

no

OpenStudy (anonymous):

SA I got is 12.6

OpenStudy (anonymous):

i must be doing something wrong

OpenStudy (anonymous):

3.14*4

OpenStudy (anonymous):

\[4\pi \text{ and} \frac{4}{3}\pi \text{ respectively} \]

OpenStudy (anonymous):

4.18

OpenStudy (anonymous):

how are your aswers supposed to be? in terms of pi?

OpenStudy (anonymous):

good question

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

the 12.6 was right

OpenStudy (anonymous):

oh so to find the volume just muliply 4/3 by pi

OpenStudy (anonymous):

1 to and positive power is 1 so you dont need to worry about anything but the pi

OpenStudy (anonymous):

ok so i got4.18

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

\[4\pi = 12.5664 \text{ and } \frac{4}{3}\pi =4.18879 \]